A 10.0 kg block is released from rest at point A. The track is frictionless except for the portion between points B and C, which has a length of 6.00m. The block travels down the track, hits a spring of force constant 2250 N/m, and compresses the spring 0.300m from its equilibrium position before coming to rest momentarily. Determine the coefficient of kinetic friction between the block and the rough surface between points B and C.

I am trying to solve this problem the following way. First find the final velocity from A to B, then I want To find The Coefficient of friction from B to C and in order to do this I believe I had to find something from C to D so I have then plug in for it in B to C.

I am looking for a baby explanation thank you.
Answer #1

initial energy Ei = mgh
final energy Ef = ½kx²
work done by friction W = μmgdcos(180) = -μmgd
W = Ef - Ei
-μmgd = ½kx² - mgh
so
μ = h/d - kx²/(2mgd)
= 3.00/6.00 - 2250*0.300²/(2*10.0*9.81*6.00)
= 0.328.

HOPE THIS HELPS!

Please thumbs up.